OpenStudy (anonymous):

What are the solutions of 3x2 = 18x? x = 0, x = 6 x = 0, x = −6 x = 3, x = 6 x = 3, x = −6

OpenStudy (anonymous):

@jigglypuff314

OpenStudy (anonymous):

Can you teach me?

jigglypuff314 (jigglypuff314):

you can divide both sides by 3 so you get x^2 = 6x and x^2 = x*x so then you can divide both sides by x

OpenStudy (anonymous):

by x you mean 18 or just x?

jigglypuff314 (jigglypuff314):

you would have divide both sides by 3 first so x^2 = 6x

OpenStudy (anonymous):

Will it be D?

jigglypuff314 (jigglypuff314):

no... whoops, sorry I had explained it wrong ignore what I said before... umm you would subtract 18x from both sides so 3x^2 - 18x = 0 then factor out a 3x using the distributive property so 3x(x - 6) = 0

OpenStudy (anonymous):

Ohhh so it would be B?

jigglypuff314 (jigglypuff314):

to find x from 3x(x - 6) = 0 you would set each part equal to 0 so 3x = 0 and x - 6 = 0

OpenStudy (anonymous):

I'm confused.

jigglypuff314 (jigglypuff314):

to made number times number = 0 true, either of the numbers has to equal 0 so to find x from 3x(x - 6) = 0 x times minus or whatever number has to equal 0 so find the values of x when 3x = 0 and x - 6 = 0

OpenStudy (anonymous):

So it's D?

jigglypuff314 (jigglypuff314):

3x = 0 and x - 6 = 0 x = 0 + 6 +6 x = 6 it would be A because x = 0 and x = 6

OpenStudy (tkhunny):

\(3x^{2} = 18x\) \(x^{2} = 6x\) \(x^{2} - 6x = 0\) \(x\cdot(x-6) = 0\) \(x = 0\;or\;x-6 = 0\) \(x = 0\;or\;x = 6\)